OpenStudy (anonymous):

Solve –4x + 7 = x + 2 x = –1 x = 1 x = –3 x = 3

OpenStudy (anonymous):

collect like terms. -4x+7=x+2 7-2=4x+x 5=5x x=5/5 x=1

OpenStudy (anonymous):

ohhh, okay thank you!

OpenStudy (anonymous):

no problem, medal thanks :)